# Density Frankl–Rödl on the Sphere

### v1

@@ Abstract
We establish a density variant of the Frankl–Rödl theorem on the sphere $\mathbb{S}^{n-1}$, which concerns avoiding pairs of vectors with a specific distance, or equivalently, a prescribed inner product. In particular, we establish lower bounds on the probability that a randomly chosen pair of such vectors lies entirely within a measurable subset $A\subseteq\mathbb{S}^{n-1}$ of sufficiently large measure. Additionally, we prove a density version of spherical avoidance problems, which generalize from pairwise avoidance to broader configurations with prescribed pairwise inner products. Our framework encompasses a class of configurations we call inductive configurations, which include simplices with any prescribed inner product $-1<r<1$. As a consequence of our density statement, we show that all inductive configurations are sphere Ramsey.
@@ 1 Introduction
The Frankl–Rödl theorem [FR87] is a foundational result in extremal combinatorics and theoretical computer science. It states that for any fixed $10<\gamma<$ 1, assuming $(1-\gamma)n$ is even, if a set $A\subseteq\{-1,1\}^{n}$ contains no pair of distinct points at Hamming distance exactly $(1-\gamma)n$, then the fractional size of $A$ must be exponentially small. That is, there exists a constant $\epsilon=\epsilon(\gamma)<1$ such that
@@
The corresponding Frankl–Rödl graph ${\mathrm{FR}_{\gamma}^{n}}$ is defined on vertex set $\{-1,1\}^{n}$ with edges between points at Hamming distance $(1-\gamma)n$, with the Frankl-Rödl theorem bounding the independence number of this graph. This theorem has found broad applications in theoretical computer science, particularly in the analysis of hardness of approximation. The graph ${\mathrm{FR}_{\gamma}^{n}}$ has been used to construct integrality gap instances for problems such as 3-Coloring [KG98, KMS98, Cha02, AG11, KOTZ14] and Vertex-Cover [KG98, Cha02, ABL02, ABLT06, GMT08, GMPT10, KOTZ14].
@@
it is natural to ask whether one can give a quantitative lower bound on this probability in terms of the density $|A|/2^{n}$. Benabbas, Hatami, and Magen [BHM12] answered this by proving a density version of the Frankl–Rödl theorem: for any $10<\alpha<$ 1, if $|A|/2^{n}\geq\alpha$, then
@@
This result is based on the reverse hypercontractivity of the Bonami–Beckner semigroup [MOR+06], and was used to prove integrality gaps for Vertex-Cover. Later Kauers et. al [KOTZ14] showed that the proof can be conducted in the sum-of-squares (SOS) proof system. In particular, they showed that for any $40<\gamma\leq 1/$ 4, the SOS/Lasserre SDP hierarchy at degree $4\left\lceil\frac{1}{4\gamma}\right\rceil$ certifies that the maximum independent set in ${\mathrm{FR}_{\gamma}^{n}}$ has fractional size $o(1)$. This implies that a degree-4 algorithm from the SOS hierarchy can certify that the ${\mathrm{FR}_{\gamma}^{n}}$ SDP integrality gap instances for 3-Coloring have chromatic number $\sigma(1)$, and that a degree-$\left\lceil 1/\gamma\right\rceil$ SOS algorithm can certify the ${\mathrm{FR}_{\gamma}^{n}}$ SDP integrality gap instances for Min-Vertex-Cover has minimum vertex cover $>1-o(1)$.
@@
For configurations $\Delta_{k}(n,r)$ with $r>0$, Castro-Silva et al. [CSdOFSV22] proved that for any $k\geq 2$, there exists $c=c(k,r)<1$ such that any $A\subseteq\mathbb{S}^{n-1}$ avoiding $\Delta_{k}(n,r)$ satisfies $\sigma(A)\leq(c+o(1))^{n}$. A configuration is sphere Ramsey if any $c$-coloring of of the sphere, there exists a monochromatic congruent copy of the configuration. Matoušek and Rödl [MR95] showed that any configuration $P$ with circumradius less than 1 is sphere Ramsey.
@@
In this paper, we show that for any $-\frac{1}{k-1}<r<1$, there exists a constant $C=C(r)$ such that
@@
for all measurable $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$. Our result generalizes the work of Castro-Silva et al. to simplices with any $-1<r<1$, and more broadly, to a class of inductive configurations defined in Section 4. As a consequence, we show that all inductive configurations are sphere Ramsey.
@@
which enables us to derive density versions of spherical avoidance for inductive configurations. In Section 2, we analyze the eigen-decomposition of $A_{t}$; in Section 3, we relate $A_{t}$ to the Poisson Markov semigroup $P_{t}$ and establish reverse hypercontractive inequalities. Finally, in Section 4, we prove our main results on density versions of the Frankl–Rödl theorem on the sphere and density spherical for inductive configurations.
@@ 2.1 Spherical Harmonics
[…]
We focus in particular on $L^{2}(\mathbb{S}^{n-1})$, the space of square-integrable functions. It admits an orthonormal basis consisting of spherical harmonics $\{S_{k,\ell}\}$, indexed by integers $k\geq 0$ (the degree) and $1\leq\ell\leq c_{k,n}$, where
@@
Each $S_{k,\ell}$ is the restriction to $\mathbb{S}^{n-1}$ of a harmonic, homogeneous polynomial of degree $k$ on $\mathbb{R}^{n}$. We denote by $\mathcal{S}_{k}\subseteq L^{2}(\mathbb{S}^{n-1})$ the finite-dimensional subspace spanned by spherical harmonics of degree $k$:
$\mathcal{S}_{k}:=\operatorname{span}\{S_{k,1},\dotsc,S_{k,c_{k,n}}\}.$
@@
[…]
where $\sigma_{x,r}$ is the uniform probability measure on the $(n-2)$-subsphere $\{y\in\mathbb{S}^{n-1}:x\cdot y=r\}$.
The operator $A_{t}$ is easily seen to commute with rotations. It is also self-adjoint:
@@
By Schur’s Lemma, any linear operator commuting with the group action must act as a scalar on each irreducible representation. Since $A_{t}$ commutes with rotations and $\mathcal{S}_{k}$ are inequivalent irreducible representations for different $k$, the restriction of $A_{t}$ to $\mathcal{S}_{k}$ must be a scalar multiple of the identity.
@@ 2.2 Eigenvalue Estimates
We now estimate the eigenvalues $\mu_{k,t}$ for different regimes of $k$:
@@
For all $0\leq k\leq n^{1/3}$, as $n\to\infty$, the eigenvalues satisfy
$|\mu_{k,t}-e^{-kt}|=O_{t}(n^{-1/3}),$
where the implicit constant depends only on $t$ but is independent of $k$.
@@
Let us first analyze the $k$-th moments of $X_{1}$. Notice that the density of $X_{1}$ is supported in $[-1,1]$ and proportional to $(1-x^{2})^{(n-4)/2}$. Since the density $X_{1}$ is an odd function, the odd moments vanishes, $\mathop{\mathbb{E}\/}[X_{1}^{k}]=0$ for odd $k$. For even $k$, there is an estimate
@@
given by [RK11] Lemma 5.5. A direct computation gives
@@
[…]
Lemma 2.3.
For all $k\geq n^{1/3}$, as $n\to\infty$, the eigenvalues satisfy
$|\mu_{k,t}|=O\left(\frac{1}{n}\right).$
Proof.
Since $\mathop{\mathbb{E}\/}[X_{1}^{2}]\leq\frac{2}{n-4}$, by Markov’s inequality,
$\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})=\mathop{\bf Pr\/}(|X_{1}|^{2}\geq\frac{1}{4})\leq\frac{8}{n-4},$
@@
$|\mu_{k,t}|$ $=\ |G_{k}(e^{-t})|$
$=\ \left|\mathop{\mathbb{E}\/}\left[\left(e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right)^{k}\right]\right|$
$\leq\ \mathop{\mathbb{E}\/}\left[\left|e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right|^{k}\right]$
$\leq\ \mathop{\mathbb{E}\/}\left[\left(e^{-2t}+(1-e^{-2t})|X_{1}|\right)^{k}\right]$
$\leq\ \mathop{\bf Pr\/}(|X_{1}|<\frac{1}{2})\left(e^{-2t}+\frac{1}{2}(1-e^{-2t})\right)^{k}+\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})$
$=\ O\left(\left(\frac{1+e^{-2t}}{2}\right)^{n^{1/3}}\right)+O(\frac{1}{n})\ =\ O(\frac{1}{n})$
@@ 3 Reverse Hypercontractivity of $A_{t}$
for $10<q<p<$ 1, whenever $t\geq\tau(p,q)$. This inequality lower-bounds the dispersion of mass under the semigroup and connects to reverse log-Sobolev inequalities [MOS13], Gaussian isoperimetry [MN15], and tail bounds in correlated settings.
@@
for all $10<q<p<$ 1.
Reverse hypercontractivity has enabled several significant applications on the hypercube. It yields anticoncentration and noise stability bounds for Boolean functions of low influence [MOO05], and dimension-free Gaussian isoperimetric inequalities [MN15].
@@ 3.1 Poisson Markov Semigroup
[…]
It is straightforward to verify that $P_{t}$ satisfies the Markov semigroup properties. Define the (unbounded) operator $L$ of the semigroup by
$Lf(x):=-\left.\frac{d}{dt}P_{t}f(x)\right|_{t=0}=\sum_{k,l}k\,\widehat{f_{k,l}}Y_{k,l}(x).$
The operator $L$ is called the generator of the Markov semigroup $P_{t}$. Then, by the semigroup property, for any $f\in L^{2}({\mathbb{S}^{n-1}})$, we have
$\frac{d}{dt}P_{t}f=-LP_{t}f=-P_{t}Lf,$
and the semigroup can be expressed formally as $P_{t}=e^{-tL}$ via the Taylor expansion of the exponential operator.
@@
$\|A_{t}f-P_{t}f\|_{2}=\left\|\sum_{k,l}(\mu_{k,t}-e^{-kt})\widehat{f_{k,l}}Y_{k,l}\right\|_{2}=\sum_{k,l}(\mu_{k,t}-e^{-kt})^{2}\left(\widehat{f_{k,l}}\right)^{2}.$
By Lemma 2.2, for $k\leq n^{1/3}$,
$(\mu_{k,t}-e^{-kt})^{2}=O_{t}(n^{-2/3}).$
By Lemma 2.3, for $k\geq n^{1/3}$,
$(\mu_{k,t}-e^{-kt})^{2}\leq\left(O\left(\frac{1}{n}\right)+e^{-tn^{1/3}}\right)^{2}=O(n^{-2}).$
Combining both regimes, we conclude
$\|A_{t}f-P_{t}f\|_{2}^{2}\leq O(n^{-2/3})\sum_{k,l}\left(\widehat{f_{k,l}}\right)^{2}=O_{t}(n^{-2/3})\|f\|_{2}^{2}\ .\qed$

### v3

@@ Abstract
We establish a density variant of the Frankl–Rödl theorem on the sphere $\mathbb{S}^{n-1}$, which concerns avoiding pairs of vectors with a specific distance, or equivalently, a prescribed inner product. In particular, we establish lower bounds on the probability that a randomly chosen pair of such vectors lies entirely within a measurable subset $A\subseteq\mathbb{S}^{n-1}$ of sufficiently large measure. Additionally, we prove a density version of spherical avoidance problems, which generalize from pairwise avoidance to broader configurations with prescribed pairwise inner products. Our framework encompasses a class of configurations we call inductive configurations, which include simplices with any prescribed inner product 1<r<1-1<r<1. As a consequence of our density statement, we show that all inductive configurations are sphere Ramsey.
@@ 1 Introduction
The Frankl–Rödl theorem [FR87] is a foundational result in extremal combinatorics and theoretical computer science. It states that for any fixed $0<\gamma<1$, assuming $(1-\gamma)n$ is even, if a set $A\subseteq\{-1,1\}^{n}$ contains no pair of distinct points at Hamming distance exactly $(1-\gamma)n$, then the fractional size of $A$ must be exponentially small. That is, there exists a constant $\epsilon=\epsilon(\gamma)<1$ such that
@@
The corresponding Frankl–Rödl graph $\mathrm{FR}_{\gamma}^{n}$} is defined on vertex set $\{-1,1\}^{n}$ with edges between points at Hamming distance $(1-\gamma)n$, with the Frankl-Rödl theorem bounding the independence number of this graph. This theorem has found broad applications in theoretical computer science, particularly in the analysis of hardness of approximation. The graph $\mathrm{FR}_{\gamma}^{n}$} has been used to construct integrality gap instances for problems such as 3-Coloring [KG98, KMS98, Cha02, AG11, KOTZ14] and Vertex-Cover [KG98, Cha02, ABLT06, GMT08, GMPT10, KOTZ14].
@@
it is natural to ask whether one can give a quantitative lower bound on this probability in terms of the density $|A|/2^{n}$. Benabbas, Hatami, and Magen [BHM12] answered this by proving a density version of the Frankl–Rödl theorem: for any $0<\alpha<1$, if $|A|/2^{n}\geq\alpha$, then
@@
This result is based on the reverse hypercontractivity of the Bonami–Beckner semigroup [MOR+06], and was used to prove integrality gaps for Vertex-Cover. Later Kauers et. al [KOTZ14] showed that the proof can be conducted in the sum-of-squares (SOS) proof system. In particular, they showed that for any $0<\gamma\leq 1/4$, the SOS/Lasserre SDP hierarchy at degree $4\left\lceil\frac{1}{4\gamma}\right\rceil$ certifies that the maximum independent set in $\mathrm{FR}_{\gamma}^{n}$} has fractional size $o(1)$. This implies that a degree-4 algorithm from the SOS hierarchy can certify that the $\mathrm{FR}_{\gamma}^{n}$} SDP integrality gap instances for 3-Coloring have chromatic number $\omega(1)$, and that a degree-$\left\lceil 1/\gamma\right\rceil$ SOS algorithm can certify the $\mathrm{FR}_{\gamma}^{n}$} SDP integrality gap instances for Min-Vertex-Cover has minimum vertex cover >1o(1)>1-o(1).
@@
For configurations $\Delta_{k}(n,r)$ with $r>0$, Castro-Silva et al. [CSdOFSV22] proved that for any $k\geq 2$, there exists $c=c(k,r)<1$ such that any $A\subseteq\mathbb{S}^{n-1}$ avoiding $\Delta_{k}(n,r)$ satisfies $\sigma(A)\leq(c+o(1))^{n}$. A configuration is sphere Ramsey if any $c$-coloring of of the sphere, there exists a monochromatic congruent copy of the configuration. Castro-Silva et al.’s result implied that all simplectic configurations $\Delta_{k}(n,r)$ with $r>0,\ k\geq 2$ are sphere Ramsey. On the other hand, Matoušek and Rödl [MR95] showed that any configuration $P$ with circumradius less than 1 is sphere Ramsey.
@@
In this paper, we show the following result:
Theorem 1.1.
Fix $k,r$ such that $\frac{1}{k-1}<r<1$, there exists constant $C=C(k,r),\ \epsilon=\epsilon(k,r)$ such that
@@
for all measurable $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega_{k,r}(n^{-\epsilon})$.
Our result generalizes the work of Castro-Silva et al. to simplices with any $\frac{1}{k-1}<r<1$, and more broadly, to a class of inductive configurations defined in Section 4 (see Theorem 4.6). In particular, we show that all inductive configurations are sphere Ramsey.
@@
which enables us to derive density versions of spherical avoidance for inductive configurations. In Section 2, we analyze the eigen-decomposition of $A_{t}$; in Section 3, we relate $A_{t}$ to the Poisson Markov semigroup $P_{t}$ and establish reverse hypercontractive inequalities. Finally, in Section 4, we prove our main results on density versions of the Frankl–Rödl theorem on the sphere and density spherical Ramsey statements for inductive configurations.
@@ 2.1 Spherical Harmonics
$\|f\|_{L^{p}(\mathbb{S}^{n-1})}:=\left(\int_{\mathbb{S}^{n-1}}|f(x)|^{p}\,d\sigma(x)\right)^{1/p},$
where $\sigma$ denotes the uniform surface measure on $\mathbb{S}^{n-1}$, normalized such that $\int_{\mathbb{S}^{n-1}}d\sigma(x)=1$. It follows that for $p\geq q>0$, we have $L^{p}(\mathbb{S}^{n-1})\subseteq L^{q}(\mathbb{S}^{n-1})$.
We focus in particular on $L^{2}(\mathbb{S}^{n-1})$, the space of square-integrable functions. It admits an orthonormal basis consisting of spherical harmonics $\{Y_{k,\ell}\}$, indexed by integers $k\geq 0$ (the degree) and $1\leq\ell\leq c_{k,n}$, where
@@
Each $Y_{k,\ell}$ is the restriction to $\mathbb{S}^{n-1}$ of a harmonic, homogeneous polynomial of degree $k$ on $\mathbb{R}^{n}$. We denote by $\mathcal{S}_{k}\subseteq L^{2}(\mathbb{S}^{n-1})$ the finite-dimensional subspace spanned by spherical harmonics of degree $k$:
$\mathcal{S}_{k}:=\operatorname{span}\{Y_{k,1},\dotsc,Y_{k,c_{k,n}}\}.$
@@
$(A_{t}f)(x):=\int_{\mathbb{S}^{n-1}}f(y)\,d\sigma_{x,e^{-t}}(y)=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}y\in{\mathbb{S}^{n-1}}\\ \ x\cdot y=e^{-t}\end{subarray}}[f(y)],$
where $\sigma_{x,r}$ is the uniform probability measure on the (n2)(n-2)-subsphere $S_{x,r}:=\{y\in\mathbb{S}^{n-1}:x\cdot y=r\}$.
The operator $A_{t}$ is easily seen to commute with rotations via
$(A_{t}Uf)(x)=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}y\in{\mathbb{S}^{n-1}}\\ \ x\cdot y=e^{-t}\end{subarray}}[f(U^{-1}y)]=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}z\in{\mathbb{S}^{n-1}}\\ \ x\cdot Uz=e^{-t}\end{subarray}}[f(z)]=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}z\in{\mathbb{S}^{n-1}}\\ \ U^{-1}x\cdot z=e^{-t}\end{subarray}}[f(z)]=(UA_{t}f)(x)$
for any $U\in\mathrm{SO}(n)$. It is also self-adjoint by
@@
By Schur’s Lemma (see e.g. [S+77]), any linear operator commuting with the group action must act as a scalar on each irreducible representation. Since $A_{t}$ commutes with rotations and $\mathcal{S}_{k}$ are inequivalent irreducible representations for different $k$, the restriction of $A_{t}$ to $\mathcal{S}_{k}$ must be a scalar multiple of the identity.
@@ 2.2 Eigenvalue Estimates
We now estimate the eigenvalues $\mu_{k,t}$.
@@
The eigenvalues $\mu_{k,t}$ satisfy
$|\mu_{k,t}-e^{-kt}|=O_{t}\left(\frac{1}{n}\right),$
as $n\to\infty$, where the implicit constant depends only on $t$ but is independent of $k$.
@@
Let us first analyze the $k$-th moments of $X_{1}$. Notice that the density of $X_{1}$ is supported in [1,1][-1,1] and proportional to $(1-x^{2})^{(n-4)/2}$. Since the density $X_{1}$ is an odd function, the odd moments vanishes, $\mathop{\mathbb{E}\/}[X_{1}^{k}]=0$ for odd $k$. For even $k$, there is an estimate
@@
given by [RK11] Lemma 5.5. We will estimate the eigenvalues $\mu_{k,t}$ for different regimes of $k$.
When $k\leq\frac{n}{10}$, a direct computation gives
@@
$\leq{}$ $\frac{2}{n-4}+\sum_{a=2}^{\left\lfloor k/2\right\rfloor}(1-e^{-2t})^{a}e^{-t(k-2a)}\left(\frac{2a}{n-4}\right)^{a}$
$\leq{}$ $\frac{2}{n-4}+\sum_{a=2}^{\left\lfloor k/2\right\rfloor}\left(\frac{2a}{n-4}\right)^{a}$
Since $\left(\frac{2a}{n-4}\right)^{a}$ is decreasing with $a$ when $2\leq a\leq\left\lfloor k/2\right\rfloor\leq\frac{n}{20}$, we can bound
$|\mu_{k,t}-e^{-kt}|\leq\frac{2}{n-4}+(\left\lfloor k/2\right\rfloor-2)\left(\frac{4}{n-4}\right)^{2}\leq\frac{2}{n-4}+\frac{n}{20}\left(\frac{4}{n-4}\right)^{2}\leq O\left(\frac{1}{n}\right).$
For the case where $k\geq\frac{n}{10}$, by Markov’s inequality
$\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})\leq\frac{\mathop{\mathbb{E}\/}\left[|X_{1}|^{2\left\lceil n/20\right\rceil}\right]}{(1/2)^{2\left\lceil n/20\right\rceil}}\leq\left(\frac{4\left\lceil n/20\right\rceil}{n-4}\right)^{n/20}\leq O\left(\frac{1}{n}\right)$
@@
$|\mu_{k,t}-e^{-kt}|$
$\leq{}$ $|G_{k}(e^{-t})|+|e^{-kt}|$
$\leq{}$ $\mathop{\mathbb{E}\/}\left[\left|e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right|^{k}\right]+|e^{-kt}|$
$\leq{}$ $\mathop{\mathbb{E}\/}\left[\left(e^{-2t}+(1-e^{-2t})|X_{1}|\right)^{k}\right]+|e^{-kt}|$
$\leq{}$ $\mathop{\bf Pr\/}(|X_{1}|<\frac{1}{2})\left(e^{-2t}+\frac{1}{2}(1-e^{-2t})\right)^{k}+\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})+|e^{-\frac{n}{10}t}|$
$\leq{}$ $\ O\left(\left(\frac{1+e^{-2t}}{2}\right)^{n/10}\right)+O(\frac{1}{n})+O_{t}(\frac{1}{n})$
$\leq{}$ $O_{t}(\frac{1}{n})$
@@ 3 Reverse Hypercontractivity of $A_{t}$
for $0<q<p<1$, whenever $t\geq\tau(p,q)$. This inequality lower-bounds the dispersion of mass under the semigroup and connects to reverse log-Sobolev inequalities [MOS13], Gaussian isoperimetry [MN15], and tail bounds in correlated settings.
@@
for all $0<q<p<1$.
Reverse hypercontractivity has enabled several significant applications on the hypercube. It was used to prove the “It Ain’t Over Till It’s Over” conjecture [Kho02] from social choice theory [MOO05], and dimension-free Gaussian isoperimetric inequalities [MN15].
@@ 3.1 Poisson Markov Semigroup
$P_{t}f(x)=\int_{{\mathbb{S}^{n-1}}}K_{e^{-t}}(x,y)f(y)\,d\sigma(y).$
It is straightforward to verify that $P_{t}$ satisfies the Markov semigroup properties.
@@
$\|A_{t}f-P_{t}f\|_{2}^{2}=\left\|\sum_{k,l}(\mu_{k,t}-e^{-kt})\widehat{f_{k,l}}Y_{k,l}\right\|_{2}^{2}=\sum_{k,l}|\mu_{k,t}-e^{-kt}|^{2}\left|\widehat{f_{k,l}}\right|^{2}.$
By Lemma 2.2,
$|\mu_{k,t}-e^{-kt}|^{2}\leq O_{t}(n^{-2})$
where we can conclude
$\|A_{t}f-P_{t}f\|_{2}^{2}\leq O_{t}\left(n^{-2}\right)\sum_{k,l}\left|\widehat{f_{k,l}}\right|^{2}=O_{t}\left(n^{-2}\right)\|f\|_{2}^{2}\ .\qed$